Using a table of critical​ t-values of the t​ distribution, find the range of values for the​ P-value for testing a claim about the mean body temperature of healthy adults for a​ left-tailed test with n=13 and test statistic t = 2.098 .

What is the range of values for the​ P-value?

A. 0.01 <P-value<0.025

B. 0.005 P-value<0.005

C.0.025<P-value<0.05

D.0.005 less than Upper P value less than 0.01
Answer #1

Solution,

Given that,

n = 13

degrees of freedom = n - 1 = 13 - 1 = 12

test statistic t = 2.098

Using t table

P ( T > 2.098 )

= 1 - P( T < 2.098 )

= 1 - 0.9711

= 0.0289

P-value = 0.0289

C.0.025<P-value<0.05